1. In a biochemical assay, a chemist needs to add 2.93 g of glucose to a reaction mixture. Calculate the volume in mL of a 1.35M glucose solution she should use for the addition.


  1. The concentrated sulfuric acid we use in the laboratory is 98.0 percent H2SO4 by mass. Calculate the molality and molarity of the acid solution. The density of the solution is 1.83g/mL
1
Expert's answer
2021-09-24T02:21:49-0400

Molar mass of glucose = 180.156

= 180.156/2.93

= 61.49ml



Consider 1 L = 1000 mL of solution

Mass of solution = volume x density

= 1000 x 1.83 = 1830 g

Mass of H2SO4 = 98.0/100 x 1830 = 1793.4 g

Moles of H2SO4 = mass/molar mass of H2SO4

= 1793.4/98.08 = 18.3 mol

Molarity = moles of H2SO4/volume of solution in L

= 18.3/1 = 18.3 M

Mass of water = 1830 - 1793.4 = 36.6 g = 0.0366 kg

Molality = moles of H2SO4/mass of water in kg

= 18.3/0.0366

= 500 mol/kg = 500m